An RLC series circuit has a 2 kΩ resistor, a 323 mH inductor, and a 36.3 LaTeX: \muμF capacitor. If this is connected to a 232 Volt power supply, what will the rms current be at 391 Hz? Express your answer in mA.
Answer #1

Here,

for the impedance, Z

Z = sqrt(R^2 + (2pi *f * L - 1/(2pi * f * C))^2)

Z = sqrt(2000^2 + (2pi * 391 * 0.323 - 1/(2pi * 391 * 36.3 *10^-6))^2)

Z = 2148 Ohm

Now, for the rms current

rms current = Vrms/Z

rms current = 232/2148

rms current = 0.108 A = 108 mA

the rms current is 108 mA